# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 4.3.2 Restriction to polytrees
We have established that the model-oriented distance for CPDAGs is inherently more complex than that for UGs or DAGs, reflecting both the increased structural complexity of the model-oriented poset and the richer graphical nature of CPDAGs themselves. We now study how the model-oriented distance may simplify if we restrict the class of CPDAGs we consider. Specifically, we focus on CPDAGs that are also polytrees, that is, CPDAGs with an acyclic skeleton. This class frequently arises in the causal inference literature (Rebane and Pearl, 1987; Sepehr and Materassi, 2019; Grüttemeier, Komusiewicz and Morawietz, 2021; Tramontano, Monod and Drton, 2022; Améndola et al., 2023)
Let $\mathfrak{G}_{\text{poly}}$ be the collection of completed partially directed graphs over the node set $V$ that are polytrees. Here, the model-oriented distance between two graphs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\text{poly}}$ is the number of neighbor steps needed to transform $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$, where each step is a move to a CPDAG that is also a polytree. Formally, let $\mathcal{L}_{\text{poly}}=(\mathfrak{G}_{\text{poly}},\preceq)$ be the poset underlying $\mathfrak{G}_{\text{poly}}$ using the same ordering relationship as the unrestricted setting (this is a subposet of $\mathcal{L}$). Then, the model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ can be characterized as follows.
Theorem 3.
@@
For any pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ in $\mathfrak{G}_{\text{poly}}$, we have $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\text{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
We prove this result in Appendix F.2. The proof proceeds by showing that $\mathcal{L}_{\text{poly}}$ is lower semimodular and then appeals to Theorem 2 to obtain the desired result.
4.3.3 Relationship to Separation Distance
There exist another distance for CPDAGs and probabilistic graphical model encoding only conditional independence statements more broadly, called separation distance (Wahl and Runge, 2025). It measures the distance between two graphs by counting the number of conditional independence statements encoded by only one of them. Similarly to our class of distances, the separation distance is a metric and model oriented, albeit on a different model interpretation. However, the separation distance framework implicitly assumes that any collection of conditional independence statements forms a valid conditional independence model. This is similar to how the SHD on CPDAGs implicitly assumes any simple graph with direct and undirected edges is a valid CPDAG (see Example 1). Conditional independence statements, however, may imply other such statements (Geiger, Verma and Pearl, 1989). In fact, determining which sets of triples correspond to a conditional independence model and which of these in turn correspond to a probabilistic graphical model is a challenging problem that has not yet been fully resolved (Sadeghi, 2017). There are two notable consequences of this idiosyncrasy. First, the separation distance between sparse graphs tends to be much larger than between dense graphs. Second, the separation distance between two graphs increases substantially when a disconnected node is added to both. We provide an example illustrating this and discuss the separation distance paradigm more thoroughly in Appendix H.2.
@@ 4.4 Causal MPDAGs
Suppose the graph class $\mathfrak{G}$ is the set of MPDAGs over vertex set $V$ with $\mathcal{M}:=\mathcal{M}_{\text{mpdag}}$ as the model map. The following result provides a graphical characterization for model containment.
@@
Let $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ be two MPDAGs. Then, $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$ if and only if for every DAG $\mathcal{D}_{1}$ represented by $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a directed subgraph of $\mathcal{D}_{2}$.
We prove Proposition 9 in Section˜A.2. Using this result, we can characterize the poset $\mathcal{L}=(P,\preceq)$ underlying MPDAGs. This poset consists of MPDAGs, ordered according to subgraph containment: graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in L$ satisfy $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if for every DAG $\mathcal{D}_{1}$ in $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$. The least element is the empty graph. The poset has a greatest element that is the fully connected undirected graph. Further, it is not meet or join semi-lattice nor is lower or upper semimodular; see Appendix E.3 for counter examples. Fig.˜3(d) shows the Hasse diagram associated with this poset. With this poset $d_{\mathcal{L}}$, we have a formal definition for model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ between two MPDAGs.
As with CPDAGs, the model‑oriented distance differs from the two variants of SHD. As a simple illustration, suppose $\mathcal{G}_{s}$ is an MPDAG among three nodes $\{v_{1},v_{2},v_{3}\}$ with two edges $v_{1}\to v_{2}$ and $v_{3}\to v_{2}$, and $\mathcal{G}_{t}$ is an MPDAG among the same node set containing two edges $v_{1}-v_{2}$ and $v_{3}-v_{2}$. Both variants of the structural Hamming distance give the value two. However, these measures implicitly assume that the graph with edge set $v_{1}-v_{2}$ and $v_{3}\to v_{2}$ is a valid MPDAG, which is incorrect — highlighting that the structural Hamming distance fails to account for model structure. In fact, the model-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ in this case is four.
Given the importance of gradedness for computing and deriving simplified characterizations of $d_{\mathcal{L}}$, we study whether the poset $\mathcal{L}$ is graded, that is, whether there exists a function $\mathrm{rank}(\cdot)$ assigning non-negative integers to graphs, such that each covering relation corresponds to a rank increase of exactly one. In previous graph classes, such a rank function was defined by the number of edges, so neighboring graphs always differed by exactly one edge. In the context of MPDAGs, however, a neighbor of a graph may contain many more edges. As an example, consider the graph $\mathcal{G}_{1}$ in Fig.˜6(a). Both graphs $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ in panels (b) and (c), respectively, satisfy $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ and $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$. The neighbor $\mathcal{G}_{2}$ contains one additional edge compared to $\mathcal{G}_{1}$, whereas $\mathcal{G}_{3}$ contains n1n-1 additional edges. Note that $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ by Proposition˜11. To see why $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$, note that any graph $\mathcal{G}$ satisfying $\mathcal{G}\preceq\mathcal{G}_{3}$, we must either direct undirected edges or drop edges. However, for $\mathcal{G}$ to also satisfy $\mathcal{G}_{1}\preceq\mathcal{G}$, none of the undirected edges in $\mathcal{G}_{3}$ can be modified in any way. Finally, dropping any subset of directed edges in $\mathcal{G}_{3}$ results in an invalid MPDAG (see Fig.˜2(b)).
This example rules out the number of edges as a valid rank function. However, it does not rule out the existence of some other valid rank function. The following proposition formally establishes that no such function exists, and thus the poset $\mathcal{L}$ is not graded.
Proposition 10.
There does not exists a function $\mathrm{rank}(\cdot)$ assigning non-negative integers to graphs in $\mathfrak{G}$, such that each covering relation corresponds to a rank increase of exactly one. Thus, the poset $\mathcal{L}$ is not graded.
We prove Proposition 10 in Appendix G.1. It constructs a pair of MPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}$ with $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$, where there are maximal chains from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ with different length. (Recall that a maximal chain in $\mathcal{L}$ is a sequence of graphs from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ in which every graph covers the proceeding one.) This difference in chain lengths contradicts the gradedness property, since every covering step increases the rank by exactly one, so any two maximal chains between the same endpoints would have to have equal length.
In addition to a lack of graded structure in $\mathcal{L}$, we do not have an explicit graphical characterization of neighbor structure. This makes it difficult to enumerate neighbors and compute $d_{\mathcal{L}}$ or the down-up and up-down distances $d_{\mathcal{L},\downarrow\uparrow}$ and $d_{\mathcal{L},\uparrow\downarrow}$. To address these challenges, we introduce a pseudo-rank function.
Definition 8 (Pseudo-rank).
For any MPDAG $\mathcal{G}\in\mathfrak{G}$, we define its pseudo-rank, denoted $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G})$, by the the number of directed edges in $\mathcal{G}$ plus twice the number of undirected edges.
The pseudo-rank function serves as a surrogate for rank in a non-graded poset (see Proposition˜11) and reflects an intuitive measure of structural complexity (see Proposition˜12).
Proposition 11 (Sufficient condition for covering).
Consider $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$. If $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{1})+1=\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{2})$, then, $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.

### v2

@@ 4.3.2 Restriction to polytrees
We have seen that the model-oriented distance for probabilistic CPDAGs does not admit a simple, closed-form expression as opposed to probabilistic UGs (Corollary˜2) and causal DAGs (Corollary˜3). This reflects the complexity of the poset underlying CPDAGs. We will use the A* algorithm to compute the distance in Section˜5. In the rest of this subsection, we show that the distance takes a much simpler form if we restrict the graph class to polytrees, namely those CPDAGs whose skeletons do not contain cycles. Due to its relative tractability, polytrees have been studied in the causal discovery literature; see, e.g., Rebane and Pearl (1987); Sepehr and Materassi (2019); Grüttemeier, Komusiewicz and Morawietz (2021); Tramontano, Monod and Drton (2022); Améndola et al. (2023).
Formally, let $\mathfrak{G}_{\mathrm{poly}}$ be the class of CPDAGs over a vertex set $\mathrm{V}$ that are polytrees and let $\mathcal{L}_{\mathrm{poly}}$ be the model-oriented poset associated with $(\mathfrak{G}_{\mathrm{poly}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}})$. The poset $\mathcal{L}_{\mathrm{poly}}$ is called the subposet of $\mathcal{L}$ induced by $\mathfrak{G}_{\mathrm{poly}}$. This subposet has the following properties, which we prove in Section˜E.3.
Theorem 6 (Polytree CPDAG poset).
Let $\mathcal{L}_{\mathrm{poly}}$ be the model-oriented poset associated with $(\mathfrak{G}_{\mathrm{poly}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}})$, where $\mathfrak{G}_{\mathrm{poly}}$ is the class of polytree CPDAGs over a finite vertex set $\mathrm{V}$. The following results hold.
  1. 1.
    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$ when $|\mathrm{V}|>2$.
  2. 2.
    For $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}_{\mathrm{poly}}$, we have $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if and only if there exist DAGs $\mathcal{D}_{1}\in[\mathcal{G}_{1}]$, $\mathcal{D}_{2}\in[\mathcal{G}_{2}]$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$.
  3. 3.
    $\mathcal{L}_{\mathrm{poly}}$ is graded with $\operatorname{\mathrm{rank}}(\cdot)$ equal to the number of edges.
  4. 4.
    $\mathcal{L}_{\mathrm{poly}}$ is lower semimodular.
  5. 5.
@@
    $d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$ holds for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\mathrm{poly}}$.
Importantly, the last result in this theorem says that we can restrict to down–up paths when finding the model-oriented distance, leading to substantial computational savings.
We conclude this subsection on CPDAGs with the following remark.
Remark 3 (Separation distance).
[…]
@@ 4.4 Causal MPDAGs
In this subsection, we consider $\mathfrak{G}$ as the set of MPDAGs over a finite vertex set $\mathrm{V}$ equipped with the model map $\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$. Recall from Section˜2.2.5 that an MPDAG $\mathcal{G}$ represents a subset of Markov equivalent (but causally distinct) DAGs that satisfy certain background knowledge; we use $\mathcal{G}]$ to denote the collection of DAGs that $\mathcal{G}$ represents. The partial order in this context is characterized graphically by the following result, which we prove in Section˜A.2; see also Lemma˜2 for comparison.
Proposition 7 (MPDAG comparability).
Consider two MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. We have $\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{s})\subseteq\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{t})$ if and only if for every DAG $\mathcal{D}_{s}\in[\mathcal{G}_{s}]$, there exists a DAG $\mathcal{D}_{t}\in[\mathcal{G}_{t}]$ such that $\mathcal{D}_{s}$ is a subgraph of $\mathcal{D}_{t}$.
Proposition˜7 is central to studying the model-oriented poset $\mathcal{L}$ of MPDAGs. Fig.˜3(d) shows the corresponding Hasse diagram when $|\mathrm{V}|=3$. Table˜2 summarizes its structural properties. In particular, the least element $\mathcal{G}_{\hat{0}}$ is the empty graph and the greatest element $\mathcal{G}_{\hat{1}}$ is the fully connected undirected graph. However, unlike UGs, DAGs and CPDAGs, the MPDAG poset does not admit the total number of edges as the rank function, as illustrated by the next example.
Example 2.
Consider the MPDAGs in Fig.˜6 over $n$ vertices. One can use Proposition˜9 to show $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$. It also holds that $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$. To see this, first observe that $\mathcal{G}_{1}\prec\mathcal{G}_{3}$ by Proposition˜7; then, note that there cannot be a graph $\mathcal{G}$ that satisfies $\mathcal{G}_{1}\prec\mathcal{G}\prec\mathcal{G}_{3}$ because $\mathcal{G}$ must result from removing one or more directed edge from $\mathcal{G}_{3}$, which would introduce a forbidden structure to the MPDAG (see Fig.˜2). Hence, despite both being a neighbor of $\mathcal{G}_{1}$, the graph $\mathcal{G}_{2}$ has one more edge than $\mathcal{G}_{1}$, while the graph $\mathcal{G}_{3}$ has n1n-1 more edges than $\mathcal{G}_{1}$.
Figure 6: MPDAGs $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}$ considered in Example˜2.
Furthermore, the following result establishes that the poset does not admit any rank function; see Section˜E.4 for its proof.
Proposition 8 (MPDAG non-gradedness).
Let $\mathfrak{G}$ be the class of MPDAGs over a finite vertex set $\mathrm{V}$ with $|\mathrm{V}|\geq 4$. The model-oriented poset of $(\mathfrak{G},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}})$ is not graded.
In addition to the lack of gradedness in $\mathcal{L}$, we do not yet have an explicit graphical characterization of the neighborhood (i.e., the covering relation); indeed, as we have seen from Example˜2, such a characterization can be technically challenging because the structures of neighboring graphs can vary greatly. Hence, it remains very difficult to enumerate neighbors, which is an essential component of any search algorithm.
Moreover, the down-up and up-down distances are less useful in the context of MPDAGs. First, $\mathcal{L}$ is neither upper nor lower semimodular (see Section˜E.4), so $d_{\mathcal{L},\downarrow\uparrow}$ and $d_{\mathcal{L},\uparrow\downarrow}$ are merely upper bounds of $d_{\mathcal{L}}$ (see Theorem˜2). Second, the lack of gradedness and semi-lattice structures (see Section˜E.4) prevents us from reducing these distances to a more computable form (see Propositions˜3 and 1).
To remedy these challenges, we introduce the following notion of pseudo-rank.
Definition 8 (MPDAG pseudo-rank).
The pseudo-rank of an MPDAG $\mathcal{G}$ is defined as
$\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}):=\text{number of directed edges}+\text{number of undirected edges}\times 2.$
This pseudo-rank serves as a surrogate for the rank in a non-graded poset and reflects an intuitive measure of structural complexity, as shown by the next two results.
Proposition 9 (MPDAG sufficient condition for covering).
For MPDAGs $\mathcal{G}_{1},\mathcal{G}_{2}$ such that $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, we have $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})\leq\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$, where the equality holds if and only if $\mathcal{G}_{1}=\mathcal{G}_{2}$. Additionally, if $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$, we have $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
Proof.
Given $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, from Proposition˜7 it follows that every $v_{1}\rightarrow v_{2}$ present in $\mathcal{G}_{1}$ corresponds to either $v_{1}\rightarrow v_{2}$ or $v_{1}-v_{2}$ in $\mathcal{G}_{2}$, and every $v_{3}-v_{4}$ present in $\mathcal{G}_{1}$ corresponds to $v_{3}-v_{4}$ in $\mathcal{G}_{2}$. The first statement then follows from definition.
Further, suppose $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$ but there exists $\mathcal{G}$ that satisfies $\mathcal{G}_{1}\prec\mathcal{G}\prec\mathcal{G}_{2}$. The skeletons of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are either identical or different by exactly one edge, which we discuss in three cases: (i) If their skeletons coincide, then $\mathcal{G}_{2}$ must have exactly one undirected edge that is directed in $\mathcal{G}_{1}$, so the intermediate $\mathcal{G}$ cannot exist. (ii) If $\operatorname{sk}(\mathcal{G}_{1})=\operatorname{sk}(\mathcal{G})$ has one fewer edge than $\operatorname{sk}(\mathcal{G}_{2})$, then the extra edge in $\mathcal{G}_{2}$ must be directed and all the other edges in $\mathcal{G}_{2}$ are identical to those in $\mathcal{G}_{1}$. Because $\mathcal{G}_{1}\prec\mathcal{G}$, $\mathcal{G}$ must have an undirected edge that corresponds to a directed edge in $\mathcal{G}_{1}$ and hence in $\mathcal{G}_{2}$, but this contradicts $\mathcal{G}\prec\mathcal{G}_{2}$. (iii) Similarly, we also have contradiction if $\operatorname{sk}(\mathcal{G})=\operatorname{sk}(\mathcal{G}_{2})$ has one more edge than $\operatorname{sk}(\mathcal{G}_{1})$. Therefore, by ruling out the intermediate $\mathcal{G}$, we conclude $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$. ∎
@@
(MPDAG model complexity) Between any MPDAG $\mathcal{G}$ and the empty graph $\mathcal{G}_{\hat{0}}$, there exists a maximal chain (i.e., of the form $\mathcal{G}_{\hat{0}}=:\mathcal{G}_{0}\lessdot\cdots\lessdot\mathcal{G}_{l}:=\mathcal{G}$) of length $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G})$.
This result is a consequence of the following lemma, which may be of independent interest.
Lemma 3 (MPDAG single-edge orientation).
Let $\mathcal{G}$ be any MPDAG with one or more undirected edges. Then, there exists an undirected edge in $\mathcal{G}$ such that by directing it and leaving all the other edges unchanged, we obtain another MPDAG $\tilde{\mathcal{G}}$.
It follows that we can turn $\mathcal{G}$ into a directed graph by orienting the undirected edges one by one, and then into the empty graph by removing the directed edges one by one; see Section˜G.1 for a constructive proof of the previous two results. Based on Proposition˜10, we define $\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ as the subset of paths between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ such that every consecutive pair of graphs on the path differ by unit pseudo-rank, namely
$\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\big\{(\mathcal{G}_{s}=:\mathcal{G}_{0},\dots,\mathcal{G}_{d}:=\mathcal{G}_{t})\in\mathrm{P}_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t}):\;\text{for $i=1,\dots,d$, }\\ |\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{i})-\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{i-1})|=1\big\}.$
In other words, $\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ only consists of paths through restricted neighborhoods, i.e., those neighboring graphs that differ by exactly one edge between a certain pair of vertices $a,b$: the difference is either aba~~~b (no edge) versus $a\rightarrow b$, or $a\rightarrow b$ versus aba-b. We then define the pseudo-distance as
$d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})\},\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$ (7)
Corollary 4.
It holds that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for every $\mathcal{G}_{s},\mathcal{G}_{t}$.